a boy throws a ball in air in such a manner that when the ball is in its max height he throws another ball. if the balls are thrown  after the time difference of 1 s ; then what will be the height attain by them:
 

Grade:11

1 Answers

Saurabh Koranglekar
3 years ago
u=0 m/s

t=1 second

time of ascend,

t=u/g

1=u/9.8

9.8 m/s\

maximum height reached, h=u^2/2g

h=(9.8)^2/2×9.8=4.9 m